# The Composition and Chemistry of Titan’s Atmosphere 111Keywords: Titan; Astrochemistry; Astrobiology; Atmospheres; Photochemistry

### v1

@@ 5.1.3 Oxygen species
Few oxygen compounds have been detected in Titan’s atmosphere: only CO, CO2 and H2O to date. Diatomic molecular oxygen, , readily found in the atmospheres of the inner planets, is absent, allowing organic chemistry to proceed to great complexity. The oxygen compounds detected appear to be attributable to an external source of oxygen (both O and OH), most likely originating from Enceladus 57, 81. Prior to the discover of the Enceladus plumes a meteoritic source was favored 506, and some meteoritic contribution may still be present 97.
@@
Photochemical models 97 predict that trace amounts of molecules such as HNO, HNCO, H2CO and CH3OH should be present in the atmosphere, and perhaps detectable at high altitudes by observatories such as ALMA. To date, few published studies have attempted to directly identify further oxygen species. The Cassini INMS team published upper limits for methanol () and acetaldehyde () of 30 ppb and 10 ppb in the ionosphere at 1100 km 219. In the stratosphere, upper limits have been determined for methanol and formaldehyde of 6 ppb and 2 ppb respectively at at 107 km, 25∘S 502 from infrared spectroscopy with Cassini CIRS.
@@ 5.2.1 Single hydrocarbon rings
Two small cyclic molecules have been definitively detected in Titan’s neutral atmosphere: cyclopropenylidene (c-C3H2) 165 and benzene (c-C6H6) 49. Other small cyclic molecules likely to exist include the saturated cycloalkanes - cyclopropane (), cyclobutane () and larger - and possibly cycloalkenes such as cyclopropene (), cyclobutene (), cyclobutadiene (), and others. Substituted rings are also possible: cyanocyclopropane ().
In the ISM, several single ring molecules have been detected. These include the small 3-carbon rings cyclopropenylidene (c-C3H2 306) and its related radical 512, and a substituted species, ethynyl cyclopropenylidene () 513. The five-sided ring cyclopentadiene (), and the six-sided rings benzyne () and benzene (c-C6H6) have also been detected 348, 513, 514.
@@ 5.2.2 Multi-ring hydrocarbons: PAHs and polyphenyls
Laboratory experimental work on haze formation by UV photolysis has shown that the inclusion of benzene in initial reagent mixtures along with N2 and CH4 leads to the formation of significantly larger molecules than when a N2/CH4 mixture is used 523. More recently, naphthalene has also be used as a starting reagent in lab tholin experiments 524 showing similar results. It should be noted that, on Titan, both benzene and naphthalene would presumably first have to form from methane, so their use in lab experiments may be considered an acceleration of a natural process that could in principle start from a pure N2/CH4 mixture and arrive at the same result over long time periods.
@@ 5.4 Radicals
For completeness, we include consideration of radical species - atoms and molecules with unpaired free electrons - even though these are highly reactive and unlikely to be found in significant numbers, or significantly deep in Titan’s neutral atmosphere. Radicals include fragments of CH4, such as CH (methylidene), CH2 (methylene), and CH3 (methyl), as well as ground state and excited nitrogen atoms formed from break up of N2, such as as N(2D) and N(4S).

### v2

@@ 5.1.3 Oxygen species
Few oxygen compounds have been detected in Titan’s atmosphere: only CO, CO and HO to date. Diatomic molecular oxygen, , readily found in the atmospheres of the inner planets, is absent, allowing organic chemistry to proceed to great complexity. The oxygen compounds detected appear to be attributable to an external source of oxygen (both O and OH), most likely originating from Enceladus 57, 81. Prior to the discover of the Enceladus plumes a meteoritic source was favored 506, and some meteoritic contribution may still be present 97.
@@
Photochemical models 97 predict that trace amounts of molecules such as HNO, HNCO, HCO and CHOH should be present in the atmosphere, and perhaps detectable at high altitudes by observatories such as ALMA. To date, few published studies have attempted to directly identify further oxygen species. The Cassini INMS team published upper limits for methanol () and acetaldehyde () of 30 ppb and 10 ppb in the ionosphere at 1100 km 219. In the stratosphere, upper limits have been determined for methanol and formaldehyde of 6 ppb and 2 ppb respectively at at 107 km, 25S 502 from infrared spectroscopy with Cassini CIRS.
@@ 5.2.1 Single hydrocarbon rings
Two small cyclic molecules have been definitively detected in Titan’s neutral atmosphere: cyclopropenylidene (c-CH) 165 and benzene (c-CH) 49. Other small cyclic molecules likely to exist include the saturated cycloalkanes - cyclopropane (), cyclobutane () and larger - and possibly cycloalkenes such as cyclopropene (), cyclobutene (), cyclobutadiene (), and others. Substituted rings are also possible: cyanocyclopropane ().
In the ISM, several single ring molecules have been detected. These include the small 3-carbon rings cyclopropenylidene (c-CH 306) and its related radical 512, and a substituted species, ethynyl cyclopropenylidene () 513. The five-sided ring cyclopentadiene (), and the six-sided rings benzyne () and benzene (c-CH) have also been detected 348, 513, 514.
@@ 5.2.2 Multi-ring hydrocarbons: PAHs and polyphenyls
Laboratory experimental work on haze formation by UV photolysis has shown that the inclusion of benzene in initial reagent mixtures along with N and CH leads to the formation of significantly larger molecules than when a N/CH mixture is used 523. More recently, naphthalene has also be used as a starting reagent in lab tholin experiments 524 showing similar results. It should be noted that, on Titan, both benzene and naphthalene would presumably first have to form from methane, so their use in lab experiments may be considered an acceleration of a natural process that could in principle start from a pure N/CH mixture and arrive at the same result over long time periods.
@@ 5.4 Radicals
For completeness, we include consideration of radical species - atoms and molecules with unpaired free electrons - even though these are highly reactive and unlikely to be found in significant numbers, or significantly deep in Titan’s neutral atmosphere. Radicals include fragments of CH, such as CH (methylidene), CH (methylene), and CH (methyl), as well as ground state and excited nitrogen atoms formed from break up of N, such as as N(D) and N(S).